In slope-intercept form, write the equation of a line that goes through point (1, 0) and (3, -4).

Answer by Cromlix(4381)

Using equation for slope (m)
m = y2 - y1/x2 - x1
m = -4 - 0/3 - 1
m = -4/2
m = -2
Using Equation y = mx + c
with coordinates (3, - 4)
y = mx + c
-4 = -2(3) + c
-4 = -6 + c
c = -4 + 6
c = 2
Equation of line =
y = -2x + 2.
Hope this helps.
:-)